# CrowdGenUI: Enhancing LLM-Based UI Widget Generation with a Crowdsourced Preference Library

### v1

@@ 5.1. UI Widget Reasoning
5.2. Integration of Crowdsourced User Preference Library
@@
5.3. UI Widget Code Generation
@@
5.4. Implementation Details
We use GPT-4o (OpenAI, 2024b) as the LLM for both widget reasoning and code generation, with Python as the programming language to implement the widgets. Occasionally, the code generated by GPT-4o may have minor bugs and might not compile correctly. We manually fix these issues to ensure all code generated by the LLM runs properly during user evaluation. This setup aligns with our focus on LLM-driven widget reasoning and code generation rather than real-time UI generation. We use this system to serve as a prototype to demonstrate how LLM reasoning, when enhanced by a crowdsourced library, can effectively guide UI widget generation according to task requirements and user preferences, particularly in the image editing domain.
6. Crowdsourcing a Preference Library for UI Widgets
In this section, we introduce how we collect user preference data to build the library that enhances LLM-driven UI widget generation. Following our previous implementation, we gather preference data in the image editing domain to illustrate the process. We present the study setup and provide an analysis of the collected data.
(a) Task description and provided UI widgets to perform the task.
(b) UI widget preference selection panel.
6.1. Platform and Participants
[…]
@@ 6.2.1. User Interface
In the crowdsourcing study, we offered each participant a Jupyter Notebook to work on eight image editing tasks, interact with several UI widgets to perform each task and provide their preferences for the widgets. Figure 4 shows the user interface for a crowdsourcing task—adjusting an image to fall colors. The interface contains the task requirement, multiple UI widgets, an image to edit, and a user preference selection panel.
@@ 6.2.2. Tasks
The crowdsourcing tasks cover three categories in the image editing domain. We introduce these categories and provide an overview of the tasks below. Some tasks may overlap across multiple categories. Our goal is to use these tasks to incorporate a range of value adjustments and data manipulation needs in image editing.
  • Continuous value adjustment: These tasks need to adjust a value across a continuous range smoothly. Examples include adjusting image lightness, saturation, hue, and color balance.
@@
The detailed task descriptions are presented in Table 1. We use these tasks to encourage exploration, such as in Tasks 1, 2, 3, 6, and 7, while also emphasizing precision, as seen in Tasks 4, 5, and 8. Two tasks are used as control tasks: image_adjust_fall_color and image_color_match. We instruct participants to ensure that the final images best meet the task requirements and use these tasks as attention checks to help identify untrustworthy responses. We randomize the order of tasks, ensuring that no particular task appears consistently in the same position to reduce order effects, such as fatigue or priming, that could influence user preferences and behaviors.
@@
Task Name Task Description
1 image_adjust_lightness Experiment with lightness settings to see how different levels affect the image.
2 image_adjust_saturation Boost the saturation to make the colors pop.
3 image_adjust_hue Experiment with different hues to find a color tone that complements the overall mood of the image.
4 image_adjust_fall_color Adjust the color settings to gradually shift the image tones, creating a warm, autumnal atmosphere. Ensure the final image is the one that looks the most autumnal to you.
5 image_color_match Change the rocket’s color to match the provided reference color. Ensure the rocket’s color in the final image best matches the reference color.
6 image_adjust_color_balance Experiment with different color balance settings to see how altering the red, green, and blue levels affects the overall color harmony of the image.
7 image_place_watermark Experiment with different watermark positions to find the perfect balance between visibility and subtlety.
8 image_place_vignette Darken the background using a vignette effect except for the human face by properly positioning the circle around the face.
@@ 6.2.3. User Preference Aspects
We use three key aspects of UI widget preferences, drawing from prior research on intuitive user interface design principles (recall Section 3). These aspects—predictability, efficiency, and explorability—reflect diverse user preferences when interacting with digital software, with the detailed descriptions below:
  • Predictability: Users can achieve results without encountering unexpected outcomes or needing to figure out how to perform the interaction.
  • Efficiency: Users can perform tasks with minimal effort.
  • Explorability: Users can explore multiple possibilities and perform tasks with high flexibility.
6.2.4. UI Widget Selection and Implementation
@@
UI widget selection needs to consider task and data manipulation requirements, such as input and output data types, and match the widget types with the specified task requirements. We choose UI widgets for the crowdsourcing tasks from the Jupyter Widgets Library (Jupyter, 2024b), which offers both basic and advanced UI widgets, covering a wide range of control needs. Specifically, for continuous value adjustment tasks, sliders [Uncaptioned image] are typically ideal to allow for fine-grained control, and text fields [Uncaptioned image] can be used for precise value input. For discrete value selection tasks, dropdown menus [Uncaptioned image] and radio buttons [Uncaptioned image][Uncaptioned image] are often effective. For tasks emphasizing visual feedback, UI widgets like color pickers [Uncaptioned image], color wheels [Uncaptioned image], direct manipulation (e.g., clicking [Uncaptioned image]), and preset buttons with preview overlays [Uncaptioned image][Uncaptioned image] allow users to see outcome previews through color or position visual cues.
In this work, we present one approach to UI widget implementation and selection, acknowledging that it does not cover all possible options. However, this method can be easily expanded by leveraging additional UI widget libraries, such as (Material, 2024; Bootstrap, 2024; Vuetify, 2024). These libraries can facilitate data manipulation and content editing in software or websites and enable the inclusion of a broader range of widget types in various applications.
6.3. Crowdsourced UI Widget Preference Library
6.3.1. Overview
@@
We distributed the crowdsourcing study to the 50 participants in three versions. Each version included all eight tasks and a subset of the three preference aspects, covering one, two, or all three aspects. This approach ensured that each participant provided their preferences for 15-18 preference selections out of the total 3 aspects $\times$× 8 tasks = 24 selections across all tasks, helping to prevent fatigue. The study yielded a UI widget preference library with 30 unique user responses for each aspect. In total, we obtained 30 responses $\times$× 3 aspects $\times$× 8 tasks = 720 user preferences. Figures 5 and 6 show the distribution of user preferences for all crowdsourcing tasks.
@@
Figure 5. User preference distribution of tasks 1-4 collected in the crowdsourcing study.
@@
Figure 6. User preference distribution of tasks 5-8 collected in the crowdsourcing study.
6.3.2. Analysis of User Preference Distribution
@@
From the crowdsourcing results of all 30 responses to each preference aspect, we observe variations in user preferences for UI widgets regarding predictability, efficiency, and explorability across the eight tasks.
For continuous value adjustment tasks, including adjusting image lightness, saturation, and hue, many users preferred preset buttons with preview overlays for their predictability and efficiency, with reasons like “The preset button’s background allows for users to figure out the lightness relatively quickly, and when clicking on the buttons, the button updates the lightness to how it looks on the buttons, allowing for users to need minimal deduction to understand what this button does”. For explorability, sliders were favored for adjusting lightness and saturation. Participants commented that “The slider allows users to slowly go through each increment and explore what effect they have on the image. They can go anywhere within the range to understand the effect of saturation on the image and precisely pinpoint which exact value of saturation they want with a slider”. A color wheel was preferred to explore hue adjustment, with comments like “The color wheel, due to it allowing you to select exactly what you want and visualizing it, allowing you to explore and experiment very easily”.
For discrete value selection tasks, including choosing predefined values for image lightness, saturation, hue, and watermark positions, preset buttons were favored over radio buttons and dropdown menus in all three preference aspects. This was primarily due to the visual cues that preset buttons provided, such as user-provided reason saying that “The preset buttons include a preview of how the saturation would look like. For example, -1.0 is fully gray, while 1.0 is a bright rainbow. This allows users who may not know what saturation means to understand what each button does; it changes how colorful the image is”.
@@
For tasks emphasizing visual feedback, preferences varied based on whether the task involved color or position adjustments. In color adjustment tasks, including changing image hue, applying fall colors, and adjusting color balance, many users preferred color wheels and color pickers for predictability and explorability, while preset buttons were preferred for efficiency, similar to continuous and discrete value adjustment tasks. However, for tasks requiring precise color adjustment, such as matching an object’s color to a given color, there was no clear consensus regarding predictability. Participants were evenly split among text fields, color wheels, color pickers, and preset buttons. The reasons for their preferences included “Since I know what my target is, just typing it gives me what I want” for the text field, “It is easy to visualize the color” for the color wheel, “The color picker would accurately allow me to pinpoint the target color green and not have a doubt that it is correct” for the color picker, and “The preset button convey closely what color the rocket will have” for the preset buttons. Regarding efficiency, most users preferred text fields, as they could quickly “enter the hex number and see the desired result”. For explorability, color wheels and color pickers were the most favored.
In position adjustment tasks, including placing a watermark or highlighting a specific area of an image, preset buttons and directly clicking on the image were preferred for predictability and efficiency, with reasons like “The presets give an indication of what to expect and wouldn’t confuse me in this task” and “Clicking is both intuitive and efficient and I know what I’$l$ get”. For explorability, users favored both clicking on the image since it was “Easy to just click in slightly different places as needed to get the area where I want it”, and using sliders because “You can totally visually discern where you want the watermarks and/or image to go to as it is continuous”.
These collected user preferences are incorporated to guide UI widget reasoning in our user evaluation of the LLM-driven UI widget generation system. We present the user evaluation details and how the crowdsourced user preferences affect reasoned widgets for user evaluation tasks in the following section.
@@ 7. User Evaluation
In this section, we introduce our user evaluation to assess our implemented system for the crowdsourced library-enhanced UI widget generation framework. We outline the research questions to evaluate in the study, describe the evaluation setup, and present the evaluation results.
@@ 7.1. Research Questions
We aim to study the following research questions in our user evaluation, focusing on the effectiveness and generalizability of the UI widget generation framework and the impact of integrating the crowdsourced UI widget preference library:
  • RQ1—Adaptivity of the framework to new, related tasks: Does the UI widget generation framework enable the generalization of crowdsourced library-enhanced widget reasoning and generation to perform tasks that differ from, but relate to, the original crowdsourcing tasks?
  • RQ2—Impact of the library on aligning generated widgets with user preferences: Can the crowdsourced UI widget preference library, which is integrated into the UI widget generation framework, help generated widgets align more closely with user preferences for predictability, efficiency, and explorability compared to those generated without a library?
  • RQ3—Effect of library size on enhancing LLM reasoning: Does the size of the crowdsourced UI widget preference library affect LLM widget reasoning effectiveness? Which size of the library provides the strongest support for LLM reasoning capabilities?
In the following subsections, we introduce the user evaluation setup designed to answer these research questions and analyze our evaluation results in alignment with them.
7.2. Platform and Participants
[…]
@@ 7.3.1. Procedure
The user evaluation invited voluntary participants, who were first presented with a consent declaration at the beginning of the study. Once participants agreed to participate, they accessed the study interface via a URL linked to our server. Before starting the tasks, they were given a briefing that outlined the study requirements. Full details of the user evaluation briefing and instructions can be found in Appendix C. Following that, they worked on three tasks and answered six questions for each task to provide their preferences for LLM-generated widgets.

### v2

@@ 4.2. UI Widget Reasoning
User context is passed to the LLM for widget reasoning. Additional input for LLM reasoning includes the crowdsourced user preference library and UI widget candidates. Crowdsourced-library ($L$) and widget-candidates ($C$) are also text data formatted as JSON. With all the input data, we instruct the LLM to use a chain-of-thought (CoT) reasoning technique to reason UI widgets. We present the CoT reasoning overview in Algorithm 1 and the detailed prompt in Appendix B.1.
The LLM first analyzes the user’s task description to identify the task requirements. Based on this analysis, it is instructed to analyze the similarity between the given task and the crowdsourcing tasks by comparing task descriptions to extract similar tasks. Once relevant tasks are identified, the LLM is further prompted to look for user-preferred widgets for each relevant task according to the preference aspect and the reasons users provided for their choices. It needs to refer to the located crowdsourced data to reason the most appropriate widgets that belong to the widget candidates. Finally, the UI widget reasoning results are formatted as JSON, including the user task’s name, the reasoned widgets for each preference aspect, and the corresponding reasoning.
Algorithm 1 Overview of CoT Reasoning
  Input: task-description ($T$), preference-aspect ($A$), crowdsourced-library ($L$), widget-candidates ($C$)
  Output: widget-reasoning
1:  for each crowdsourced-task in $L$ do
2:     relevant-tasks $\leftarrow$ similarity ($T$, crowdsourced-task) {Identify relevant tasks}
3:     for each aspect in $A$ do
@@
5:        return  ui-widgets $\in$ $C$Return matching widgets}
6:     end for
7:  end for
@@
4.3. UI Widget Implementation
Once the reasoned widgets are obtained, the LLM is further used to generate the Python code to implement these widgets. During this process, we provide an example code, which covers the implementation of all the widget candidates, as guidance for the LLM to reference during code generation. Our experiment shows that including example code significantly reduces bugs in the LLM-generated code. More importantly, the example code enables the LLM to successfully generate widget implementations that go beyond the default functionality of the Jupyter Widgets Library. For instance, a color wheel widget requires additional use of the Interactive Canvas in Jupyter (Jupyter, 2024a) and the implementation of mouse interactions on the wheel to map clicks to the corresponding colors. In addition to specific widget implementations, the example code also offers a structure to organize and display widgets in a clear, structured format within Jupyter Notebook for user interaction. The detailed prompt and example code for UI widget implementation can be found in Appendix B.2.
4.4. System Implementation Rationale
Task domain: We implement the system in the image editing domain because it offers generalizability across a diverse range of content editing domains. We carefully design image editing tasks that involve various data manipulation needs, including continuous and discrete value adjustments and tasks emphasizing color and position visual feedback (Section 5.3.2). These tasks require both precise control and iterative exploration, which are essential for content editing requirements in general. Moreover, the interactions employed in image editing, characterized by real-time visual feedback, trial and error, and fine-tuning content attributes, are directly transferable to other digital content editing applications such as video editing, data visualization, and graphic design. Consequently, image editing tasks address fundamental data manipulation needs that apply to broader digital content editing practices.
User interface: The current implementation uses Jupyter Notebook to build the user interface as it can handle all the data manipulation requirements for our tasks. Although this implementation does not cover other UI possibilities, such as web or professional software, the framework is extendable to accommodate these UIs by updating the widget candidates to web or software widget libraries and generating code in the corresponding programming languages for widget implementation.
UI widget candidates: The current implementation focuses on the Jupyter Widgets Library widgets rather than open-ended options. This allows practical feasibility since implementing an infinite variety of widgets can be unmanageable. Also, these widgets are strategically chosen as foundational elements that map to universal interaction principles. For instance, sliders and text fields for continuous adjustments can extend to more granular controls like range sliders or numeric steppers, and dropdowns and radio buttons for discrete selections can evolve into multi-select interfaces or dynamic filters. Focusing on these core widgets can allow scalability by integration with external libraries (e.g., Angular (Material, 2024), React Bootstrap (Bootstrap, 2024), or Vuetify (Vuetify, 2024)) to expand into advanced UI components in future iterations.
LLM selection: We use GPT-4o (OpenAI, 2024b) for UI widget reasoning and implementation due to its advanced reasoning capabilities, robust code generation performance, and strong alignment with natural language instructions, which are critical for dynamically generating UIs based on task descriptions and user preferences.
5. User Preference Crowdsourcing
In this section, we introduce our collected user preference data to build the library for UI generation. We present why we use crowdsourcing to collect this data, followed by the crowdsourcing setup and an analysis of the collected data.
@@ 5. User Preference Crowdsourcing
This figure shows an example of the crowdsourcing user interface consisting of three components. From top to bottom, the components are the task description, the UI widget interaction panel to perform the task, and the user preference selection panel. The task description specifies the goal that the user is expected to achieve. Below it, the UI widget interaction panel shows multiple widgets on the left and an image to edit on the right. At the bottom, the preference selection panel allows users to select their preferred widgets and write their reasons for the three preference aspects.
5.1. Motivation
User preference is part of user feedback and plays a crucial role in UI design and UI widget selection. The user feedback data is traditionally collected through labor-intensive methods like interviews, usability tests, A/B tests, and feedback forms (Stone et al., 2005; Wilson, 2013). Given that LLMs are trained on large-scale human data, they provide a promising solution to incorporating user feedback, such as user preferences, into dynamically generated UIs. However, without dedicated user preference datasets, current LLMs cannot fully address this need (recall Section 2.3). To fill this gap, we propose crowdsourcing a preference library to enhance LLM-based UI widget generation with user preference data, motivated by the effectiveness of adopting crowdsourcing for user feedback collection to assist various design areas in previous research endeavors (Xu et al., 2014; Hossain and Kauranen, 2015; Xu et al., 2015; Luther et al., 2015; Yu et al., 2016; Foong et al., 2017; Lee et al., 2018; Oppenlaender et al., 2020).
5.2. Participants
The crowdsourcing took place on Prolific (Prolific, 2024). We recruited participants with Python programming experience only to ensure they were familiar with running the provided Python code in Jupyter Notebook. During the study, they did not write code but just executed our provided code to interact with UI widgets and work on their tasks on desktops using a mouse and keyboard. We did not require participants to have image editing or UI programming experience, as we aim to collect regular users’ preference data and make generated UIs align with end users’ preferences.
We acknowledge that this participant recruitment may limit our sample by excluding users who are unfamiliar with Python. However, these participants can still meaningfully contribute to general users’ UI preferences, as programming expertise is independent of UI usage and preference for non-programming tasks.
According to prior research on user testing for UX design (Tomczak et al., 2023; University, 2025; Interviews, 2025), 20–30 participants are generally sufficient for evaluating simple to intermediate tasks, while 50 participants are recommended for more complex tasks. Following these established guidelines, we invited 50 participants to the crowdsourcing study (Age: 29.86 $\pm$ 7.83, Male: 37, Female: 12, Non-binary: 1), and compensated them 10 USD for the 30-minute study.
@@ 5.3.1. User Interface
During the crowdsourcing study, each participant had a Jupyter Notebook to work on eight image editing tasks, interact with multiple UI widgets to perform each task, and provide their preferences for the widgets. Figure 4 shows the UI for a crowdsourcing task—adjusting an image to fall colors. The interface contains the task requirement, multiple UI widgets, an image to edit, and a user preference selection panel.
@@ 5.3.2. Tasks
The crowdsourcing tasks cover three categories in the image editing domain. Some tasks may overlap across multiple categories. Our goal is to use these tasks to incorporate a range of value adjustments and data manipulation needs.
Continuous value adjustment: These tasks need to adjust a value across a continuous range. Examples include adjusting image lightness, saturation, hue, and color balance.
@@
The detailed task descriptions are presented in Table 1. We use these tasks to encourage exploration, such as in Tasks 1, 2, 3, 6, and 7, while also emphasizing precision, as seen in Tasks 4, 5, and 8. Tasks 4 and 5 are used as control tasks for attention checks to help us identify untrustworthy responses, where we instruct participants to ensure that the final images best meet the task requirements. We randomize the order of tasks, ensuring that no particular task appears consistently in the same position to reduce order effects that could influence user behaviors.
@@
[…]
This table consists of two columns: task name and task description. In each row, the number, name, and description of each crowdsourcing task are detailed. Task Name Task Description 1 image_adjust_lightness Experiment with lightness settings to see how different levels affect the image. 2 image_adjust_saturation Boost the saturation to make the colors pop. 3 image_adjust_hue Experiment with different hues to find a color tone that complements the overall mood of the image. 4 image_adjust_fall_color Adjust the color settings to gradually shift the image tones, creating a warm, autumnal atmosphere. Ensure the final image is the one that looks the most autumnal to you. 5 image_color_match Change the rocket’s color to match the provided reference color. Ensure the rocket’s color in the final image best matches the reference color. 6 image_adjust_color_balance Experiment with different color balance settings to see how altering the red, green, and blue levels affects the overall color harmony of the image. 7 image_place_watermark Experiment with different watermark positions to find the perfect balance between visibility and subtlety. 8 image_place_vignette Darken the background using a vignette effect except for the human face by properly positioning the circle around the face.
@@ 5.3.3. User Preference Aspects
We use predictability, efficiency, and explorability as user preference aspects, drawing from prior research on intuitive user interface design principles (recall Section 4.1.2). (1) Predictability emphasizes that users can achieve results without encountering unexpected outcomes or needing to figure out how to perform the interaction. (2) Efficiency highlights that users should be allowed to perform tasks with minimal effort. (3) Explorability underscores that users can explore multiple possibilities and perform tasks with high flexibility.
5.3.4. UI Widget Candidates
@@
As presented in Section 4.4, we aim to select UI widgets to cover various data manipulation requirements and align these candidates with the specified task requirements. To this end, the Jupyter Widgets Library (Jupyter, 2024b) offers both basic and advanced UI widgets, including a wide range of control needs. Specifically, for continuous value adjustment tasks, sliders are selected to allow for fine-grained control, and text fields can be used for precise value input; for discrete value selection tasks, dropdown menus and radio buttons are often effective; for tasks emphasizing visual feedback, UI widgets like color pickers, direct manipulation (e.g., clicking), and preset buttons with preview overlays are picked to show outcome previews through color or position cues.
5.4. Crowdsourced UI Preference Library
5.4.1. Statistics Overview
@@
We distributed the crowdsourcing study to our participants in three versions. Each version included all eight tasks and a subset of the three preference aspects, covering one, two, or all three aspects. This approach ensured that each participant provided their preferences for 15-18 preference selections out of the total 3 aspects $\times$ 8 tasks = 24 selections across all tasks, helping to prevent fatigue. The study yielded a UI widget preference library with 30 unique user responses for each aspect. In total, we obtained 30 responses $\times$ 3 aspects $\times$ 8 tasks = 720 user preferences. Figures 5 and 6 show the distribution of collected user preferences.
@@
Figure 5. User preference distribution of tasks 1-4 (task set 1) collected in the crowdsourcing study.
[…]
This figure consists of four rows, showing the crowdsourcing results of tasks 1 to 4 in each row. Each row has three columns, depicting the user preference distribution of the three preference aspects in each column as bar charts.
@@
Figure 6. User preference distribution of tasks 5-8 (task set 2) collected in the crowdsourcing study.
[…]
This figure consists of four rows, showing the crowdsourcing results of tasks 5 to 8 in each row. Each row has three columns, depicting the user preference distribution of the three preference aspects in each column as bar charts.
5.4.2. Analysis of User Preferences
@@
We observe variations in user preferences for UI widgets regarding predictability, efficiency, and explorability across the eight tasks.
Continuous value adjustment: Users preferred preset buttons with preview overlays for their predictability and efficiency. For explorability, sliders were favored for adjusting lightness and saturation, while a color wheel was preferred to explore hue adjustment.
Discrete value selection: Preset buttons were generally favored over radio buttons and dropdown menus in all three preference aspects, primarily due to the visual cues provided.
@@
Tasks emphasizing visual feedback: User preferences varied based on whether the task involved color or position adjustments. (1) In color adjustment tasks emphasizing exploration, many users preferred color wheels and color pickers for predictability and explorability, while preset buttons were selected for efficiency, similar to continuous and discrete value adjustment tasks. (2) However, for tasks requiring precise color adjustment, such as matching an object’s color to a given color, there was no clear consensus regarding predictability. Participants were evenly split among text fields, color wheels, color pickers, and preset buttons. Regarding efficiency, most users preferred text fields. For explorability, color wheels and color pickers were the most favored. (3) On the other hand, in position adjustment tasks, preset buttons and directly clicking on the image were preferred for predictability and efficiency, while for explorability, users favored both clicking on the image and using sliders.
@@ 6. User Evaluation
In this section, we introduce the user evaluation to assess our implemented system for the CrowdGenUI framework. We outline the research questions to evaluate in the study, describe the evaluation setup, and present the evaluation results.
@@ 6.1. Research Questions
We aim to study the following research questions in our user evaluation, focusing on the effectiveness and generalizability of the proposed framework and the impact of integrating the user preference library on UI generation:
RQ1—Adaptivity of the framework to new, related tasks: Does the framework enable the generalization of crowdsourced library-enhanced widget reasoning and generation to perform tasks that differ from, but relate to, the original crowdsourcing tasks?
RQ2—Impact of the library on aligning generated widgets with user preferences: Can the crowdsourced UI widget preference library help LLM-generated widgets align more closely with user preferences for predictability, efficiency, and explorability compared to those generated without a library?
RQ3—Effect of library size on enhancing LLM reasoning: Does the size of the crowdsourced UI widget preference library affect LLM widget reasoning effectiveness? Which size of the library provides the strongest support for reasoning capabilities?
In the following subsections, we explain the rationale of these RQs, introduce the user evaluation setup designed to answer them, and analyze our evaluation results in alignment with them.
6.2. Participants
We conducted the user evaluation on Prolific (Prolific, 2024) and employed the same recruitment criteria as those in the crowdsourcing study to ensure consistency in participant background. The participants worked on the user evaluation tasks on desktops using a mouse and keyboard. In total, we recruited 78 participants (Age: 29.81 $\pm$ 9.09, Male: 56, Female: 19, Non-binary: 3) and compensated each participant 10 USD for the 30-minute study.
@@ 6.3.1. Procedure
The user evaluation invited voluntary participants, who were first presented with a consent declaration at the beginning of the study. Once participants agreed to participate, they accessed the study interface via a URL linked to our server. Before starting the tasks, they were given a briefing that outlined the study requirements. Full details of the user evaluation briefing and instructions can be found in Appendix C. Following that, they worked on three tasks and answered six questions for each task to provide their feedback on LLM-generated widgets.